8.02 ESG Independent Study

Unit 13: Magnetic Properties of Matter

At this point, we are prepared to investigate one of the oldest and most profound questions in physics: How can magnets stick shopping lists onto refrigerators? We know that magnetic fields are produced by moving charges or currents and we know that magnetics fields exert forces on moving charges and currents, but permanent magnets seem to exert forces on each other in the absence of conventional currents. This unit considers another way of regarding currents in order to explain this earliest known magnetic phenomenon, and in doing so introduces a reformulation of Ampère’s Law which accommodates permanent magnets.

Magnetic properties of materials are of concern in transformers, propagation of radiation (Units 15 & 16), certain kinds of new-fangled cookware, and superconductors (coming soon to a temperature near you).

Objectives: After completing this unit, you should be able to understand and explain the phenomenon of diamagnetism, paramagnetism and ferromagnetism, use Ampère’s Law as applicable to magnetic materials, and explain why refrigerators stick to magnets.

Suggested Procedure:

1. Reread section 28.8 in UP11. Those interested in a more in-depth approach or better preparation for the test are strongly advised to read Purcell. Suggested problems include 41, 42, 43, 82.

or,

2. Read chapter eleven in Purcell. Section 3 leans a bit on the math, but be sure you understand the qualitative result, shown graphically in figures 11.7 & 11.8. Suggested problems include (pp. 442–450) #s 2, 7, 11, 12, 16, & 23. Problem 27, while interesting and important, is beyond the scope of this course (it involves quantum mechanics and vector potentials), but don’t let that stop you if you’re so inclined.

3. Take a unit test.